2500 kg pickup accelerates from 0 to 100 km/hr in 10 s along level road. What would be corresponding time when pulling 500 kg trailer? Assume constant acceleration and neglect all retarding forces.
1
Expert's answer
2012-08-14T10:03:34-0400
When pickup accelerates from 0 to 100 km/h (=100*1000/3600 m/s = 27.(7) m/s) in 10 s the acceleration of a pickup is

a1 = 27.(7)m/s / 10s = 2.(7) m/s².

The corresponding force is

F = m*a1 = 2500kg * 2.(7) m/s² = 6944.(4) N.

The acceleration when pulling 500 kg trailer will be

a2 = 6944.(4)N/(2500kg + 500kg) = 2.3(148) m/s².

So, the pickup pulling 500 kg trailer will accelerate from 0 to 100 km/h in

t = 27.(7)m/s / 2.3(148)m/s² = 12 s.